OpenStudy (anonymous):

Solve. 4(a + 2) = 14 – 2(3 – 2a)

jigglypuff314 (jigglypuff314):

4(a + 2) = 14 – 2(3 – 2a) distribute 4(a)+4(2)=14+(-2)(3)+(-2)(-2) then simplify 4a + 8 = 14 - 6 + 4 4a + 8 = 12 then subtract 8 from both sides -8 -8 4a = 4 then divide both sides by 4 /4 /4 a = 4/4 so a =